OpenStudy (anonymous):

Factor the difference of 2 cubes: 54x^3-250y^3

OpenStudy (debbieg):

Do you know the rule for factoring difference of cubes?

OpenStudy (anonymous):

a^3-b^3= (a-b)(a^2+ab+b^2)

OpenStudy (debbieg):

\[\huge a^3 – b^3 = (a – b)(a^2 + ab + b^2)\]

OpenStudy (debbieg):

yes. So just factor out the common factor - there is a GCF of 2 - and then you should have a difference of cubes to just plug and chug in that formula.

OpenStudy (anonymous):

I hate math and I'm not good at it at all. I just don't have a "math brain." i know the GCF is 2... but I dont know what to do after that.. 2(27x^3-125y^3) ??

OpenStudy (ybarrap):

Note: \(\large 54=3^3\times 2\) and \(\large 250=5^3\times 2\). So there are some nice cubes in there with a common factor.

OpenStudy (anonymous):

2x(3x^2-5x^2) ?

OpenStudy (debbieg):

If you multiply that 2 back through what you did, do you get what you started with?

OpenStudy (debbieg):

"factoring out" just means you take that common factor (2 is the GCF because it's the biggest number that is a factor of both 54 and 250) and "divide it out" of each term, so that it comes out side the ( )'s.

OpenStudy (debbieg):

\[54x^3-250y^3=2(?x^3-?y^3)\]

OpenStudy (ybarrap):

Close: \(\Large 2((3x)^3 - (5y)^3)\), then set \(a=3x\) and \(b=5x\) in our formula above

OpenStudy (anonymous):

(3x-5x)(6x+8x+25x) ?

OpenStudy (ybarrap):

You got it! I should have said set \( a=3x\) and \(b=5y \) in our formula: \((3x-5y)(9x^2+15xy+25y^2) \), which is as simple as we can get. That's it!

OpenStudy (anonymous):

ok, thank you @ybarrap wouldnt it be (3x-5y)(9x+15xy+25x)

OpenStudy (anonymous):

i mean (3x-5y)(9x+15x+25y)

OpenStudy (anonymous):

ah, i keep missing one: (3x-5y)(9x+15xy+25y)

OpenStudy (ybarrap):

The \(9x\) should be \(9x^2 \) because when we set \(a=3x\) and then square it, we get \((3x)^2=9x^2\), same thing for \(25x\): \(b^2=(5y)^2=25y^2 \) and in the middle, instead of 15x, you should have for \(ab=(3x)(5y)=15xy\)

OpenStudy (anonymous):

its saying that's not correct @ybarrap

OpenStudy (ybarrap):

did you multiply the whole thing times 2? Remember this was our common factor between 54 and 250 that allowed us to set a=3x and b=5y: \(2(3x-5y)(9x^2+15xy+25y^2)\)

OpenStudy (anonymous):

no. so its... (6x-10y)(18x^2+30xy+50y^2) ?

OpenStudy (ybarrap):

\(2(3x-5y)(9x^2+15xy+25y^2)=(6x-10y)(9x^2+15xy+25y^2)\) Remember, for example, 2*(5)(6) = (2*5)(6) or (5)(2*6) not (2*5)(2*6), which can easily verify with a calculator

OpenStudy (ybarrap):

Does this make sense?

OpenStudy (anonymous):

yes, it said that was wrong too. it was looking for 2(3x-5y)(9x^2+15xy+25y^2)

OpenStudy (ybarrap):

That's exactly what I have above

OpenStudy (anonymous):

i put in the (6x-10y) one. Thank you though